How can I combine 1/2logc+9logd and express them as a single logarithm?
Answer by lwsshak3(6761)

How can I combine 1/2logc+9logd and express them as a single logarithm?
1/2logc+9logd
1/2logc=logc^(1/2)
9logd=logd^9
addition in logs indicates multiplication in operations.
Placing under single log:
1/2logc+9logd=log[c^(1/2)*(d^9)]